# mini-vec2vec: Scaling Universal Geometry Alignment with Linear Transformations

### v2

@@ Abstract
We build upon vec2vec, a procedure designed to align text embedding spaces without parallel data. vec2vec finds a near-perfect alignment, but it is expensive and unstable. We present mini-vec2vec, a simple and efficient alternative that is highly robust and incurs a much lower computational cost. Moreover, the learned mapping is a linear transformation. Our method consists of three main stages: a tentative matching of pseudo-parallel embedding vectors, transformation fitting, and iterative refinement. Our linear alternative exceeds the original instantiation of vec2vec by orders of magnitude in efficiency, while matching or exceeding their results. The method’s stability and interpretable algorithmic steps facilitate scaling and unlock new opportunities for adoption in new domains and fields.
Algorithm 1 Unsupervised Embedding Alignment
1:Embedding matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$, $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$
2:Number of clusters $c$, neighbors $k$, iterations $T$, smoothing parameter $\alpha$, runs $s$, clusters $c^{\prime}$, neighbors $k^{\prime}$, sample size $n_{s}$
3:Linear transformation matrix $\mathbf{W}$
4:Normalize embeddings: $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}\leftarrow\text{center and normalize to unit sphere}$
5:Create pseudo-pairs: $\mathcal{C}\leftarrow\operatorname{AnchorAlignment}(\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B},c,k,s)$ $\triangleright$ See Alg. 2
6:Estimate initial mapping: $\mathbf{W}\leftarrow\text{Procrustes analysis on }\mathcal{C}$
7:Refine-1: Apply matching-based refinement on $\mathbf{W}$ $\triangleright$ See Alg. 3
8:Refine-2: Apply clustering-based refinement on $\mathbf{W}$ $\triangleright$ See Alg. 4
9:return $\mathbf{W}$
@@ 1 Introduction
Representation learning has revolutionized natural language processing by producing embedding models that capture rich semantic structure in high-dimensional vector spaces. These learned representations encode complex linguistic relationships, enabling downstream tasks such as semantic similarity and information retrieval. However, different models, training procedures, or even different random initializations of the same architecture can produce embedding spaces that, while semantically meaningful, are not directly comparable due to arbitrary rotations, reflections, and translations in their coordinate systems.
@@
The problem of aligning embedding spaces without parallel supervision – that is, without access to pairs of embeddings that represent identical semantic content – has emerged as a fundamental challenge with far-reaching implications. Successful unsupervised alignment has multiple implications across many domains: multilingual natural language processing (Conneau et al., 2018; Zhang et al., 2017), cross-modal alignment between text and image representations (Maniparambil et al., 2024), and privacy concerns (Song and Raghunathan, 2020; Morris et al., 2023).
The recent vec2vec method (Jha et al., 2025) demonstrated that unsupervised alignment is indeed possible using CycleGAN (Zhu et al., 2020), an adversarial training framework. Their approach is theoretically grounded in the Platonic Representation Hypothesis (Huh et al., 2024), which posits that well-trained representations, regardless of their specific architectures or training procedures, tend to converge toward geometrically similar spaces. Under this hypothesis, semantic relationships are preserved across models in terms of relative distances and angles, suggesting that alignment should be achievable through geometric transformations.
While vec2vec has proven effective, adversarial training frameworks bring inherent challenges that limit their practical applicability. GANs are notoriously computationally intensive, requiring careful hyperparameter tuning and often exhibiting training instability characterized by mode collapse, oscillatory behavior, or failure to converge (Saxena and Cao, 2023; Goodfellow, 2017). These methods typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
@@
Our point of departure is still universal geometry, but we pursue a different approach. We are motivated by the remarkable success of supervised alignment approaches, particularly neural network stitching methods (Lenc and Vedaldi, 2015; Bansal et al., 2021), which have demonstrated that simple affine transformations often suffice to align representations. Unfortunately, stitching requires access to pairs of matching representations. Our key methodological insight is that with universal geometry, structural similarities should be detectable and exploitable, even without adversarial training. Specifically, the relative arrangement of semantic clusters, the preservation of neighborhood structures, and the overall geometric organization of concepts should provide sufficient signal for alignment, even in the complete absence of parallel supervision. To this end, we utilize the Relative Representation framework of Moschella et al. (2023), which allows us to operate in a universal space agnostic of arbitrary coordinate systems.
@@
Our contribution is mini-vec2vec, a simple and robust pipeline with competitive alignment performance offering several key advantages over adversarial approaches: computational efficiency (running on commodity CPU hardware), training stability (deterministic components with controllable stochastic elements), interpretability (each step has a clear geometric interpretation), and sample efficiency (effective with substantially fewer training examples). Through comprehensive experiments, we demonstrate that this simple approach not only matches but often exceeds the performance of the more complex adversarial method while requiring orders of magnitude less computational resources.
@@ 2 Preliminaries
The hypothesis that neural networks converge to geometrically similar representations has been explored across multiple domains and scales. Early work by Olah (2015) revealed striking structural regularities in representations of different convolutional networks, showing distances between object representations are roughly the same regardless of architecture. These findings suggested that the geometric organization of learned features might be more universal than previously assumed. The Platonic Representation Hypothesis, formally articulated by Huh et al. (2024), represents the most comprehensive theoretical framework for understanding universal geometry. The hypothesis posits that well-trained representations converge toward a shared geometric structure that reflects the underlying structure of the data domain.
@@
Moschella et al. (2023) operationalized the observations of Olah (2015) and proposed relative representations as a mathematically principled framework for comparing latent spaces. Their key insight was that while absolute coordinates may vary arbitrarily across models, relative similarity structures – captured through pairwise distances to a shared set of anchor data points – remain remarkably stable. This work provided both a theoretical foundation and empirical evidence for the existence of model-agnostic geometric properties.
Given a set of anchor points $\{\mathbf{a}_{i}\}_{i=1}^{k}$ in an embedding space, the relative representation of any point $\mathbf{x}$ is defined as:
$\mathbf{r}(\mathbf{x})=[f(\mathbf{x},\mathbf{a}_{1}),f(\mathbf{x},\mathbf{a}_{2}),\ldots,f(\mathbf{x},\mathbf{a}_{k})]^{T},$
where $f(\cdot,\cdot)$ is a similarity function (typically cosine similarity). Two representations from different models will have close relative representations if they represent similar objects.
Algorithm 2 Algorithm for creating pseudo-pairs using noisy anchors
1:Normalized embeddings $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}$, clusters $c$, neighbors $k$, runs $s$
2:Set of pseudo-parallel pairs $\mathcal{C}$
3:Initialize collections: $\mathcal{R}_{A}\leftarrow[],\mathcal{R}_{B}\leftarrow[]$
4:for $i=1$ to $s$ do $\triangleright$ Multiple runs for robustness
5:  Cluster both spaces: $\{\mathbf{c}^{A}_{i,j}\}\leftarrow\text{k-means}(\hat{\mathbf{X}}_{A})$, $\{\mathbf{c}^{B}_{i,j}\}\leftarrow\text{k-means}(\hat{\mathbf{X}}_{B})$
6:  Compute similarities: $\mathbf{S}^{A},\mathbf{S}^{B}\leftarrow\text{cosine similarities between centroids}$
7:  Find correspondence: $\mathbf{P}^{*}_{i}\leftarrow\text{solve QAP}(\mathbf{S}^{A},\mathbf{S}^{B})$ using multiple 2-OPT runs
8:  Build relative representations: $\mathbf{r}^{A}_{i},\mathbf{r}^{B}_{i}\leftarrow\text{cosine similarities to anchors}$
9:  Store: append $\mathbf{r}^{A}_{i},\mathbf{r}^{B}_{i}$ to $\mathcal{R}_{A},\mathcal{R}_{B}$
10:end for
11:Concatenate: $\mathbf{r}^{A}\leftarrow\text{concatenate all runs in }\mathcal{R}_{A}$
12:Concatenate: $\mathbf{r}^{B}\leftarrow\text{concatenate all runs in }\mathcal{R}_{B}$
[…]
14:Create pseudo-pairs: $\mathcal{C}\leftarrow\left\{(\mathbf{x}^{A}_{i},\frac{1}{k}\sum_{j}\mathbf{x}^{B}_{j}):\mathbf{r}^{B}_{j}\in\mathcal{N}_{k}(\mathbf{r}^{A}_{i})\right\}$
15:return $\mathcal{C}$
@@ 3 Problem Formulation
We follow the setup laid out in vec2vec. We assume access to two pools of text embeddings, each from a different encoder, and our task is to align their embedding spaces without paired examples. Similar to vec2vec, we learn a parameterized mapping (neural network or linear transformation) rather than a one-to-one matching between data points. This approach offers key advantages:
@@
  • Out-of-distribution capability: The parameterized nature enables generalization beyond the training data
@@
The universal geometry hypothesis indicates that such a task can be possible: “good enough” models learn equivalent geometries where cosine similarities remain approximately invariant across embedding spaces. These geometric similarities enable the identification of this alignment.
Let $\mathcal{X}_{A}=\{\mathbf{x}^{A}_{i}\}_{i=1}^{n_{A}}$ and $\mathcal{X}_{B}=\{\mathbf{x}^{B}_{j}\}_{j=1}^{n_{B}}$ denote two sets of $d$-dimensional embeddings, where $\mathbf{x}^{A}_{i},\mathbf{x}^{B}_{j}\in\mathbb{R}^{d}$. These embeddings represent semantic content from potentially different models, architectures, or training procedures, but we assume they encode similar semantic information from the same underlying distribution $\mathcal{D}$. Unlike more traditional setups, there does not exist a pairing between data points, i.e., there is no overlap between the sentences themselves.
@@
We organize these embeddings into matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$ and $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$, where each row corresponds to a single embedding vector. The fundamental challenge is that while both spaces may contain similar semantic structures, they are related by an unknown transformation. Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(\mathbf{x}^{A}_{i})\approx\mathbf{x}^{B}_{j}$ when $\mathbf{x}^{A}_{i}$ and $\mathbf{x}^{B}_{j}$ represent similar semantic content. More generally, the setup allows for learning a pair of functions $f_{1},f_{2}$ sending embeddings to a third space, such that $f_{1}(\mathbf{x}^{A}_{i})\approx f_{2}(\mathbf{x}^{B}_{j})$. We will not take advantage of this flexibility in the paper. [^1] Of course, $f_{1}(x)=f_{2}(x)=\text{const}$ would be a trivial solution, without additional constraints. So, similarity needs to be evaluated in comparison to other vectors in the dataset.
Algorithm 3 Refine-1: Matching-Based Refinement
1:Embedding matrices $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}$, initial transformation $\mathbf{W}$
2:Iterations $T$, smoothing parameter $\alpha$, neighbors $k^{\prime}$, sample size $n_{s}$
3:Refined transformation matrix $\mathbf{W}$ after matching-based refinement
4:for $t=1$ to $T$ do
5:  Sample embeddings: $\mathbf{X}_{\text{sample}}\leftarrow\text{random sample of }n_{s}\text{ rows from }\hat{\mathbf{X}}_{A}$
6:  Transform samples: $\mathbf{X}_{\text{transformed}}\leftarrow\mathbf{X}_{\text{sample}}\mathbf{W}$

### v3

@@ Abstract
We build upon vec2vec, a procedure designed to align embedding spaces without parallel data. While it usually finds a near-perfect alignment, it is expensive and unstable. We present mini-vec2vec, a simple and efficient alternative that is highly robust and incurs a much lower computational cost. Moreover, the transformation that mini-vec2vec learns is linear. The method consists of three main stages: a tentative matching of pseudo-parallel embedding vectors, transformation fitting, and iterative refinement. Our linear alternative exceeds the original instantiation of vec2vec by orders of magnitude in efficiency, while matching or exceeding its alignment performance. The method’s stability and interpretable steps facilitate scaling and unlock new opportunities for adoption in new domains and fields.
@@ 1 Introduction
Sentence embeddings capture rich semantic structure in high-dimensional vector spaces. They encode complex linguistic relationships and enable downstream tasks, such as semantic similarity and information retrieval. However, different architectures, training procedures, and even different models with the same architecture produce embedding spaces that cannot be compared directly, because they use a different coordinate system, a different “internal language”. The problem of aligning embedding spaces without parallel supervision – that is, without pairs of matching encodings of the same data – has emerged as a fundamental challenge with far-reaching implications. Successful unsupervised alignment has multiple implications across many domains: multilingual natural language processing (conneau2017word; zhang-etal-2017-adversarial), cross-modal alignment between text and image representations (maniparambil2024; schnaus2025), and privacy concerns (song2020informationleakage; text2vec).
The recent vec2vec method (jha2025vec2vec) has demonstrated that unsupervised alignment is indeed possible using CycleGAN (zhu2017cyclegan). Their approach is theoretically grounded in a principle called the Universal Geometry or the Platonic Representation Hypothesis (huh2024platonic), which posits that well-trained representations, regardless of their specific architectures or training procedures, tend to converge toward geometrically similar spaces, even if their coordinate systems differ. This means that there is a way to rotate the coordinate systems so they “understand” each other – and vec2vec does precisely that; it identifies a rotation that makes the coordinate systems aligned. While vec2vec has proven effective, adversarial training frameworks like CycleGAN bring inherent challenges that limit their practical applicability. Training them is computationally intensive, requires careful hyperparameter tuning, and is subject to training instabilities like mode collapse, oscillatory behavior, or failure to converge (saxena2023gan_survey; goodfellow2017nips2016). Moreover, they typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
@@
Our point of departure is still Universal Geometry, but we take a different approach. We are motivated by the remarkable success of supervised alignment approaches, particularly neural network stitching methods (lenc2014stitching; bansal2021revisiting), which have demonstrated that simple affine transformations often suffice to align representations. But unfortunately, stitching assumes access to pairs of matching representations. Motivated by the Platonic Representation Hypothesis, we hypothesize that we can nonetheless uncover this linear alignment without even needing adversarial training. Specifically, we hypothesize that the similarity in the geometric structure – the relative arrangement of semantic clusters, the preservation of neighborhood structures, and the overall geometric organization of concepts – should provide sufficient signal for alignment without parallel supervision and without need for adversarial methods. To this end, we utilize the Relative Representations framework of moschella2022relative, which allows us to operate in a universal space agnostic of arbitrary coordinate systems. We then use it as a stepping stone to produce a noisy alignment that we refine with further steps. [^1] In Section 6, we discuss a variant that sidesteps the use of relative representations, but we tested it far less.
@@
mini-vec2vec is a simple and robust pipeline with competitive alignment performance offering several key advantages over adversarial approaches: computational efficiency (very fast and running on commodity CPU hardware), training stability (good performance and robustness to hyperparameters), interpretability (each step has a clear geometric interpretation), and sample efficiency (effective with substantially fewer training examples). Through comprehensive experiments, we demonstrate that this simple approach not only matches but often exceeds the performance of the more complex adversarial method while requiring orders of magnitude less computational resources.
@@ 2 Preliminaries
The hypothesis that neural networks converge to geometrically similar representations has been explored across multiple domains and scales. Early work by colah2015visualizing has revealed striking structural similarities in representations of different convolutional networks, [^2] Olah briefly refers to even earlier ideas (e.g., a concept called “meta-tSNE” by Hinton), but we failed to locate them. suggesting that the geometric organization of learned features might be more universal than previously assumed. The Platonic Representation Hypothesis, formally articulated by huh2024platonic, represents the most comprehensive theoretical framework for understanding Universal Geometry. The hypothesis posits that well-trained representations converge toward a shared geometric structure that reflects the underlying structure of the data domain. Having the same geometry implies that the distance between objects should be independent of the models themselves. For example, if one wants to compute the semantic similarity of “I love your dog” and “I love my cat” using cosine similarity between their embeddings, the answer should be roughly the same regardless of the underlying embedding model. Intuitively, this empirical observation stands to reason, as semantic similarity originates from the semantic meaning that comes from the data itself. Below, we discuss the framework of moschella2022relative that extends this observation into a powerful tool.
Relative Representations
@@
moschella2022relative have operationalized the observations of colah2015visualizing and proposed to use relative representations as a mathematically principled framework for comparing latent spaces. While absolute coordinates vary arbitrarily across models, we know that relative information, such as similarity between two objects, is (approximately) agnostic to the choice of embedding model. They propose to use a shared set of anchor points that other data points will be compared to as a way to ground the position of data points in a model-agnostic way.
Denote the space of data points by $\mathcal{Z}$ (e.g., sentences, images, or any other modality) and the embedding function by $E:\mathcal{X}\rightarrow\mathbb{R}^{d}$. Given a set of anchor data points $\{\mathbf{a}_{i}\}_{i=1}^{k}\subseteq\mathcal{Z}$, the relative representation of $\mathbf{z}\in\mathcal{Z}$ under $E$ is:
$\mathbf{r}_{E}(\mathbf{z})=[f(E(\mathbf{z)},E(\mathbf{a}_{1}));f(E(\mathbf{z}),E(\mathbf{a}_{2}));\ldots;f(E(\mathbf{z}),E(\mathbf{a}_{k}))],$
where $f(\cdot,\cdot)$ is a similarity function (typically cosine similarity) and ; represents concatenation.
The power of relative representations is that, as shown by colah2015visualizing and moschella2022relative, for two encoders $E_{1},E_{2}$, the relative representations of the same data point is roughly the same, i.e., $\mathbf{r}_{E_{1}}(\mathbf{z})\approx\mathbf{r}_{E_{2}}(\mathbf{z})$. In other words, the function $\mathbf{r}_{E}:\mathcal{X}\rightarrow\mathbb{R}^{D}$ is itself an embedding that maps a data point into $k$-dimensional space in a way that is approximately independent of $E$ (for a sufficiently well-trained $E$) and only depends on the choice of anchor points.
@@ 3 Problem Formulation
We follow the setup laid out in vec2vec. We assume access to two pools of text embeddings, each from a different encoder, and our task is to align their embedding spaces without paired examples. Similar to vec2vec, we learn a parameterized mapping (neural network or linear transformation) rather than a one-to-one matching between data points. This approach offers several advantages:
@@
  • Out-of-distribution capability: The parameterized nature can (potentially) enable generalization beyond the training distribution
@@
The Universal Geometry hypothesis indicates that such a task can be possible: “good enough” models learn equivalent geometries – relative similarities between data points remain approximately the same regardless of the embedding model.
Let $\mathcal{X}_{A}=\{\mathbf{x}^{A}_{i}\}_{i=1}^{n_{A}}$ and $\mathcal{X}_{B}=\{\mathbf{x}^{B}_{j}\}_{j=1}^{n_{B}}$ denote two sets of $d$-dimensional embeddings, where $\mathbf{x}^{A}_{i},\mathbf{x}^{B}_{j}\in\mathbb{R}^{d}$. The sentences come from the same underlying distribution $\mathcal{D}$, but – unlike most other works – there is no overlap between the sentences themselves – only the distribution is shared. We organize these embeddings into matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$ and $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$, where each row corresponds to a single embedding vector. The fundamental challenge is that while both spaces may contain similar geometric structures, they are related by an unknown transformation.
@@
Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(E_{A}(\mathbf{z}))\approx E_{B}(\mathbf{z})$ where $\mathbf{z}$ represents a datapoint $\mathbf{z}\in\mathcal{Z}$. More generally, the setup described in jha2025vec2vec allows for learning a pair of functions $f_{A},f_{B}$ sending embeddings to a third space, such that $f_{A}(E_{A}(\mathbf{z}))\approx f_{B}(E_{B}(\mathbf{z}))$. We will not take advantage of this flexibility in this paper. [^3] It also complicates the analysis of the results. For example, $f_{A}=f_{B}=\text{const}$ would be a trivial solution if we don’t impose additional constraints. Similarity then needs to be assessed more carefully.
4 Methodology
4.1 Motivation
In this work, we restrict our attention to linear transformations, i.e., $f(\mathbf{x})=\mathbf{W}\mathbf{x}$ where $\mathbf{W}\in\mathbb{R}^{d\times d}$. Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}\mathbf{W}^{T}=\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$. This is motivated by the Universal Geometry hypothesis – we expect cosine similarities to be preserved across embedding spaces. Later on, we will relax this assumption and ultimately the matrix will be an exponentially-smoothed average of orthogonal matrices, enforcing orthogonality softly.
Our linear approach, mini-vec2vec, addresses key limitations of the current approach: GANs are notoriously unstable and hard to train, demand extensive computational resources and training time. They often require tedious hyperparameter tuning and sometimes multiple re-runs to converge to good solutions. On the other hand, we know that in supervised cases (model stitching), there exists an affine mapping between the latent spaces, suggesting the linear case is realizable – the question is how to find it efficiently. mini-vec2vec achieves that, and it has several appealing properties: it requires significantly less data than vec2vec (60k vs. 1 million samples), runs on a CPU rather than requiring extensive GPU resources, and provides more stable training dynamics.
Our method resembles the approach described in hoshen-wolf-2018-non for word embeddings, proceeding in three main stages: approximate matching, transformation fitting, and iterative refinement. This paradigm is very powerful, and we adopt a similar approach. Unfortunately, text embeddings are somewhat more complicated than word embeddings, partly because there is no small exhaustive set of objects (words) that need to be matched. Moreover, the very concept of non-overlapping datasets is somewhat contrived for word embeddings, precisely because the set of possible objects is closed and relatively small. Therefore, word embedding alignment is rarely considered in the non-overlapping datasets setup (even in multilingual matching, words mostly have a ground-truth translation). We find that applying the algorithm from hoshen-wolf-2018-non naively to our use case doesn’t work. The approximate matching stage, in particular, requires more delicate treatment in our case. The iterative refinement step is also improved by a slightly more subtle approach.
def mini_vec2vec(X_A, X_B):
X_A, X_B = normalize(X_A), normalize(X_B)
C = anchor_pairs(X_A, X_B)
W = procrustes(C)
W = refine_1(X_A, X_B, W)
W = refine_2(X_A, X_B, W)
return W
def anchor_pairs(X_A, X_B):
R_A_list, R_B_list = [], []
for i in range(num_runs):
C_A = k_means(X_A, num_clusters)
C_B = k_means(X_B, num_clusters)
S_A = similarity_function(C_A, C_A)
S_B = similarity_function(C_B, C_B)
pi = solve_QAP(S_A, S_B)
C_B_permuted = permute(C_B, pi)
R_A_list.append(similarity_function(X_A, C_A))
R_B_list.append(similarity_function(X_B, C_B_permuted))
R_A, R_B = concatenate(R_A_list), concatenate(R_B_list)
neighbors = knn(R_A, R_B, num_neighbors=k)
X_matched = average_nearest_neighbors(X_B, neighbors)
return X_A, X_matched
Listing 1: Pseudo-code for mini-vec2vec (refinement algorithms deferred to Listing LABEL:lst:refine)
We will now describe our algorithm in detail. Pseudo-code is provided in Listing LABEL:lst:full_algo.
4.2 Preprocessing
We first normalize the embeddings in both spaces. We center each set around its respective mean and normalize the embeddings to the unit hypersphere:
4.3 Approximate Matching
To establish structural correspondences, we employ an anchor discovery procedure based on multiple runs of clustering and matching. We will first describe a single run, and then explain how we robustify it by ensembling multiple runs:
Centroid Matching
The reason we choose to align cluster centroids is that they roughly represent the central landmarks in the dataset. Since the data comes from a shared distribution $\mathcal{D}$, we expect clusters to be similar overall, and all that remains to do is to find the alignment between them. Of course, clusters are not very consistent even with exactly the same data, so the approach needs some tolerance to errors (hence the robustification step below).
We perform k-means clustering on both embedding spaces independently to obtain cluster centroids $\{\mathbf{c}^{A}_{j}\}$ and $\{\mathbf{c}^{B}_{j}\}$ where $j=1,\ldots,k$. We then compute similarity matrices between the centroids independently in each embedding space:
We then find the optimal matching between cluster centroids from different embedding spaces by solving the Quadratic Assignment Problem (QAP):
$\pi^{*}=\arg\max_{\pi\in\Pi_{k}}\;\sum_{i=1}^{k}\sum_{j=1}^{k}S^{A}_{ij}\,S^{B}_{\pi(i)\,\pi(j)}$
where $\pi:\{1,2,...,k\}\rightarrow\{1,2,...,k\}$ is a permutation from the set $\Pi_{k}$ of all permutations on $k$ elements. Due to the small number of cluster centroids, this takes only a few seconds.
Relative Representations
Once we have identified an alignment between $A\text{-}$ and $B\text{-centroids}$, we use the aligned centroids as anchors. We represent each embedding through its relationships to anchor points. For embedding $\mathbf{x}^{A}_{i}$, we construct:
$\mathbf{r}^{A}_{i}=[\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{1}),\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{2}),\ldots,\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{k})]$
Similarly, for $\mathbf{x}^{B}_{i}$, we compute (note that we need to re-arrange the indices with $\pi$):